# Language Detection by Means of the Minkowski Norm: Identification Through Character Bigrams and Frequency Analysis

### v1

@@ Abstract
The debate surrounding language identification has gained renewed attention in recent years, especially with the rapid evolution of AI-powered language models. However, the non-AI-based approaches to language identification have been overshadowed. This research explores a mathematical implementation of an algorithm for language determinism by leveraging monograms and bigrams frequency rankings derived from established linguistic research. The datasets used comprise texts varying in length, historical period, and genre, including short stories, fairy tales, and poems. Despite these variations, the method achieves over 80% accuracy on texts shorter than 150 characters and reaches 100% accuracy for longer texts and older writings. These results demonstrate that classical frequency-based approaches remain effective and scalable alternatives to AI-driven models for language detection.
@@ 1 Introduction
Language identification is a fundamental task in natural language processing (NLP) [1], with broad applications across machine learning [2], information retrieval [3], and text classification [4]. Approaches to this task generally fall into two categories: AI-based methods [5], such as deep learning and neural networks, and non-AI-based, statistical methods [6]. While AI-based techniques dominate current research, they often require large training datasets, substantial computational resources, and complex model management.
In contrast, statistical methods leveraging character frequencies, diacritics, and bigrams provide lightweight and accessible alternatives. Prior studies have investigated n-gram classification and character frequency analysis [7], but there is limited research on diacritic pattern utilization, and to our knowledge, no work has combined these features systematically. This paper proposes a novel approach that integrates these non-AI techniques to improve language identification accuracy, particularly for texts of at least 150 characters, while maintaining robustness in challenging scenarios.
@@ 2.3 Non-AI-Based Approaches
Early search engines used heuristic and statistical methods, including n-gram models and URL analysis, to infer language [3].
Before the rise of AI, statistical models such as n-gram analysis were the primary tools for language identification [4]. These methods compare the frequency distributions of character sequences (monograms, bigrams, etc.) in a text to language-specific profiles [7]. Frequency-based approaches remain attractive for their simplicity, interpretability, and low computational overhead. However, their accuracy can be challenged by short texts, code-switching (i.e., alternating between two or more languages or language varieties within a single conversation or even within a sentence), or closely related languages [14, 15, 16, 17, 18, 19, 20].
@@ 2.5 Datasets
We evaluate our approach using four datasets: the Fairy Tales Data Set (FTDS), the Language Determinism Data Set (LDDS), the OPUS Multilingual Parallel Corpus, and the ROST dataset.
@@
Source: povesti-pentru-copii.com
@@
Source: GitHub repository
@@
Source: OPUS website
@@ 3 Methodology and Algorithm Development
This section details the iterative development of our non-AI language identification algorithm, emphasizing the mathematical and empirical rationale for each design choice. We will use C++ as the programming language for our implementation because it offers faster execution and better performance than other high-level languages like Python, especially when processing large datasets. This choice supports an efficient implementation capable of handling high data volumes.
@@ 4.6 Comparative Performance Analysis
To quantitatively validate the effectiveness of our integrated approach, we conducted rigorous testing across four datasets:
@@ 4.7 Limitations and Discussion
The primary remaining limitation is the restricted language coverage (currently 5 languages). Expanding to logographic languages (e.g., Chinese) or highly inflectional languages (e.g., Finnish) would require additional feature engineering.
@@ 4.8 Summary of Key Findings
5 Conclusion
This research has established a robust, non-AI methodology for language detection using character-level statistical features. By leveraging monogram and bigram frequency distributions with diacritic pattern analysis, our approach achieves high accuracy (up to 100% for texts $>150$ characters) while avoiding the computational overhead of neural methods. The algorithm’s design prioritizes interpretability and extensibility, providing a transparent alternative to black-box AI solutions. C++ was used in this project because it offers faster execution and better performance than other high-level languages like Python, especially when processing large datasets. This choice supports an efficient implementation capable of handling high data volumes.

### v2

@@ Abstract
The debate surrounding language identification has gained renewed attention in recent years, especially with the rapid evolution of AI-powered language models. However, the non-AI-based approaches to language identification have been overshadowed. This research explores a mathematical implementation of an algorithm for language determinism by leveraging monograms and bigrams frequency rankings derived from established linguistic research. The datasets used comprise texts varying in length, historical period, and genre, including short stories, fairy tales, and poems. Despite these variations, the method achieves over 80% accuracy on texts shorter than 150 characters and reaches 100% accuracy for longer texts. These results demonstrate that classical frequency-based approaches remain effective and scalable alternatives to AI-driven models for language detection.
@@ 1 Introduction
Language identification is a fundamental task in natural language processing (NLP) [8], with broad applications across machine learning [6], information retrieval [5], and text classification [3]. Approaches to this task generally fall into two categories: AI-based methods, such as deep learning and neural networks [2], and non-AI-based, statistical methods [9]. While AI-based techniques dominate current research, they often require large training datasets, substantial computational resources, and complex model management.
In contrast, statistical methods leveraging character frequencies, diacritics, and bigrams provide lightweight and accessible alternatives. Prior studies have investigated n-gram classification and character frequency analysis [11]. However, there is limited research on diacritic pattern utilization, and to our knowledge, no work has combined these features systematically. This paper proposes a novel approach that integrates these non-AI techniques to improve language identification accuracy, particularly for texts of at least 150 characters, while maintaining robustness in challenging scenarios.
@@ 2.3 Non-AI-Based Approaches
Early search engines used heuristic and statistical methods, including n-gram models and URL analysis, to infer language [5]. Before the rise of AI, statistical models such as n-gram analysis were the primary tools for language identification [3]. These methods compare the frequency distributions of character sequences (monograms, bigrams, etc.) in a text to language-specific profiles [11]. Frequency-based approaches remain attractive for their simplicity, interpretability, and low computational overhead. However, their accuracy can be challenged by short texts, code-switching (i.e., alternating between two or more languages or language varieties within a single conversation or even within a sentence), or closely related languages [4, 7, 10, 16, 17, 18, 20].
@@ 2.5 Datasets
As we already stated in Section 1, we evaluate our approach using four datasets, which are described in detail in Table 1.
@@
Source: https://www.povesti-pentru-copii.com
@@
Source: https://github.com/PogaceanPaul/DataSetLanguage/tree/main
@@
Source: http://opus.nlpl.eu/
@@ 3 Methodology and Algorithm Development
This section details the iterative development of our non-AI language identification algorithm, emphasizing the mathematical and empirical rationale for each design choice. We will use C++ as the programming language for our implementation because it offers faster execution and better performance than other high-level languages like Python [19], especially when processing large datasets. This choice supports an efficient implementation capable of handling high data volumes.
@@ 4.6 Comparative Performance Analysis
To quantitatively validate the effectiveness of our integrated approach, we conducted rigorous testing across all four datasets:
@@ 4.7 Limitations and Discussion
The primary remaining limitation is the restricted language coverage (currently 5 languages). Expanding to logographic languages (e.g., Chinese) or highly inflectional languages (e.g., Finnish) may require additional feature engineering.
@@ 4.8 Summary of Key Findings
5 Conclusions
This research has established a robust, non-AI methodology for language detection using character-level statistical features. By leveraging monogram and bigram frequency distributions with diacritic pattern analysis, our approach achieves high accuracy (up to 100% for texts longer than 150 characters) while avoiding the computational overhead of neural methods. The algorithm’s design prioritizes interpretability and extensibility, providing a transparent alternative to black-box AI solutions. C++ was used in this project because it offers faster execution and better performance than other high-level languages like Python, especially when processing large datasets. This choice supports an efficient implementation capable of handling high data volumes.